OpenStudy (anonymous):

dy/dx of xy=(x^2)+(y^2)

OpenStudy (starwars18):

x'y+xy' = 2xx' + 2yy'

OpenStudy (anonymous):

The answer posted for this problem is (2x-y)/(x-2y)

OpenStudy (phi):

can you do the implicit differentiation ? of for example \[ \frac{d}{dy}(xy) \]

OpenStudy (landon34):

just use http://www.geteasysolution.com/

OpenStudy (anonymous):

Its for a final i just wanted to understand it

OpenStudy (anonymous):

I got the right answer i just wanted to know how that one guy got x'y+xy' = 2xx' + 2yy'

OpenStudy (anonymous):

Since i was able to manipulate it to the correct answer

OpenStudy (phi):

by the product rule \[ \frac{d}{dx}(xy) = x \frac{d}{dx}y + y \frac{d}{dx}x \\ = x\frac{dy}{dx}+ y \]

OpenStudy (anonymous):

Thanks!

OpenStudy (phi):

and the right-hand side is \[ \frac{d}{dx}x^2 + \frac{d}{dx}y^2 \\= 2 x + 2y \frac{dy}{dx}\]

OpenStudy (anonymous):

I think i just over thought it

OpenStudy (phi):

ok, and to finish it (and using y' for dy/dx) btw x' in this case would mean dx/dx = 1 xy' + y = 2x + 2y y' x y' - 2y y' = 2x -y (x - 2y) y' = 2x -y and \[ y' = \frac{2x-y}{x-2y} \]